Solve the system of linear equations by graphing.

?=−2?+3

−2?=?+1
Answer #1

Solve. x-into y-into - 2x + 3 y = - 28+3. - 2x = y + 1 (95 0) (0,3) (1o) (0, -1) boraph Ac Parallel lines lines are parallel.